PLEASE HELP ME ANSWER THESE QUESTIONS!!D Question 10 1 pts Use the following information for questions 10-14. Below are results of a test using an acid-free bacillus smear to screen for active pulmonary tuberculosis in patients who were diagnosed with the disease and in controls that did not have the disease No Tuberculosis 50 750 Total 240 760 Tuberculosis Positive Screening Result 190 Negative Screening Result Total 10 200 800 1,000 Calculate the sensitivity. A 95.0% O B. 92.0% ° C. 88.0% D. 78.0% Question 11
Answer #1

10)Sensitivity

Given poso

positive 190 in 200

190÷200×100=95%

11)Specificity

Given 750 are in negative screening out of 800

750÷800×100=93.75% (93.8%)

12)Positive predictive value:

Given 190 case are positive

total 240

190÷200×100=79.1 or 79.2%

13)Negative predictive value is

Given

750 are negative out if 760

750÷760×100=98.68% that is 98.7%